# ProSGNeRF: Progressive Dynamic Neural Scene Graph with Frequency Modulated Auto-Encoder in Urban Scenes

### v2

@@ 3.2 Background Model
Our local scene graph contains a background representation node, shown at the top of Fig. 1, which models the background and static objects. The static background module function maps a coordinate $x$ to its density and color:
${[\sigma(\boldsymbol{x}),\boldsymbol{z}]}$ $=F_{\theta_{bkg,1}}\left(\gamma_{x}(\boldsymbol{x})\right)$ (3)
@@
3.3 Frequency Auto-Encoder
All existing NeRF methods have difficulties in multi-dynamic objects in urban scenes. The original NeRF method assumes the density $\sigma$ is static, but both the density $\sigma(\boldsymbol{x})$ and color $\boldsymbol{c}(\boldsymbol{x})$ depend on time (and video). SUDS [21] uses an optical flow method to differentiate dynamic objects from the background. However, their method cannot model dynamic objects effectively, resulting in blurred object portions and numerous artifacts in the rendered images. Furthermore, most of the dynamic objects only appear in a few observations. This setting starkly contrasts the majority of existing works that observe multiple views of the same object. To address this challenging problem, we propose a frequency auto-encoder in dynamic objects node to learn a normalized, object-centric high-level feature that describes and disentangles shape, appearance, and pose.
@@ 3.3 Frequency Auto-Encoder
where $S_{o}$ is the inverse length of the bounding box size $s_{o}=[L_{o},H_{o},W_{o}]$ with $So=diag(1/so)$. The scaling factor allows the network to learn size-independent representation. We share the weights of the network for the objects with a similar appearance and shape.
@@
Shape and Appearance Encoder. For a input image $I$, we use the segment anything model and 3D object detector to get the 2D mask $M$ and 3D bounding box $\beta$. We encode the input images $I$ depicting a given object of interest into a shape code $\mathcal{S}_{\theta}$ and an appearance code $\mathcal{A}_{\theta}$ via a neural network. Inspired by [15], we design a two-head output architecture with the CNN feature extractor for generating the shape code $l_{s}$ and appearance code $l_{a}$, respectively.
[…]
@@
Frequency Modulation. In real-world urban scenes, dynamic objects only appear in a few observations, which significantly differs from the settings of the original NeRF method. In sparse view settings, the Neural radiance field is prone to overfitting to these 2D images without explaining 3D geometry in a multi-view consistent way, leading to reconstruction failure. Prior work [25] has already demonstrated that this issue is presumably caused by the high-frequency feature of the input.
@@
To this end, we further improve the feature-extracting network with a frequency regularization method. The original auto-encoder addressed the sparsity issue by utilizing a CNN network to encode the high-frequency features, which leads to the overfitting of high-frequency features. In this work, we use a linearly increasing frequency mask $\boldsymbol{\alpha}$ to regulate the visible frequency spectrum of the position embedding function:
@@
Figure 3: Qualitative results on reconstruction and novel scene arrangements of a scene from the KITTI dataset [8] for NeRF [14], NSG [16], SUDS [21] and our method. From left to right, these images correspond to different timesteps captured in the dynamic scene. We place the PSNR values of each scene in the bottom right corner. NeRF and SUDS are limited in their ability to properly represent the dynamic parts of the scene. In contrast, our neural scene graph method achieves high-quality view synthesis results for both image reconstruction and novel scene synthesis, regardless of scene dynamics.
@@
KITTI [8] KITTI - 75% KITTI - 50% KITTI - 25%
@@
NeRF [14] 18.54 0.563 0.431 18.50 0.554 0.438 18.12 0.547 0.447 17.81 0.535 0.468
NeRF$+$Time 23.18 0.647 0.205 20.85 0.602 0.272 21.34 0.615 0.255 18.55 0.566 0.267
NSG [16] 25.34 0.744 0.192 23.26 0.673 0.204 22.74 0.659 0.216 21.13 0.632 0.225
SUDS [21] 26.95 0.805 0.153 22.77 0.797 0.095 23.12 0.821 0.095 20.76 0.747 0.117
Ours 27.37 0.857 0.153 24.35 0.823 0.086 24.75 0.834 0.082 22.35 0.792 0.095
Table 1: Image reconstruction and novel view synthesis results on the complete KITTI [8] 0006 sequence. We compare our method with others on PSNR, SSIM, and LPIPS metrics. For PSNR and SSIM, higher is better; for LPIPS, lower is better. Our method generates higher-quality rendering images both in image reconstruction and novel view synthesis.
All the additional information related to scene graph dynamic objects nodes: the coordinate $\boldsymbol{x}$, direction $\theta\phi$, and object pose is embedded by our frequency encoding before they are put into the auto-encoder. Our frequency auto-encoder significantly improves the scene representation of dynamic objects and achieves great performance in image reconstruction, novel view synthesis, and object editing.
@@ 3.4 Scene Graph Rendering
Ray-Box Sampling. For each ray $r$ in the dynamic object node, we use the ray-box sampling method. We check the rays that intersect with the dynamic object nodes and translate them to the object coordinate space. Then, we apply the efficient AABB intersection method [12] to compute all ray-box entrance and exit points $(t_{o1},t_{oN})$. We stratified sample points $\boldsymbol{x}_{o}$ between the entrance and exit points:
@@
Volume Rendering. The sample points along the rays can be represented as:$\left\{t_{i}\right\}_{i=1}^{N_{s}+m_{j}N_{d}}$. $m_{o}$ represents the number of dynamic nodes that the ray intersects along its path. The color $c(x_{i})$ and volumetric density $\sigma(x_{i})$ of each intersection point is predicted from the respective network in the static background node $F_{\theta_{bkg}}$ or dynamic objects node $F_{\theta_{o}}$. Then, we define the termination probability and color as:
@@ 3.5 Progressive Scene Graph Joint Learning
For the large-scale dynamic scene, we divide the entire scene into multi-local scene graphs. Whenever the camera moves to the bound of the current scene graph, we dynamically initialize a new neural scene graph and stop optimizing the previous ones. In each local scene graph, we use color loss to optimize the representation model for the background and dynamic object node:
@@
where $\Pi()$ is the projection from 3D point $P_{i}$ to the depth of 2D pixel. $R_{i}$ is the set of rays with a valid depth projection value. Inspired by [5], we also use ray distribution loss to minimize the KL divergence between the rendered ray distribution $h(t)$ and the noisy depth distribution:
$\mathcal{L}_{\sigma}=\sum_{i}logh_{i}\mathrm{exp}(-\frac{(t_{i}-\hat{\boldsymbol{D}_{i}})^{2}}{2\hat{\sigma}_{i}^{2}})\delta_{i}$ (12)
@@
Methods Image Reconstruction Novel View Synthesis
KITTI* [8] KITTI* - 75% KITTI* - 50% KITTI* - 25%
[…]
NeRF [14] 16.98 0.509 0.524 16.81 0.509 0.524 16.64 0.504 0.535 16.55 0.498 0.544
NeRF + Time 24.18 0.677 0.237 22.05 0.612 0.262 21.90 0.602 0.239 20.17 0.591 0.253
NSG [16] 26.66 0.857 0.071 25.25 0.801 0.089 24.81 0.786 0.105 22.31 0.752 0.134
SUDS [21] 28.31 0.857 0.065 25.21 0.834 0.144 24.29 0.821 0.156 19.73 0.758 0.217
Ours 30.31 0.891 0.055 26.75 0.853 0.084 25.55 0.841 0.089 23.35 0.802 0.109
Table 2: We report the PSNR, SSIM, and LPIPS metrics results on the subsequent scenes from KITTI 0006 [8], following the setting of NSG and SUDS. Our approach outperforms prior methods of reconstruction and novel view synthesis across all evaluation metrics for image quality.
@@
[…]

### v3

@@ 3.2 Background Model
[…]
$N_{l}$ $=\left\lfloor N_{\min}\cdot b^{l}\right\rfloor,$ (3)
$b$ $=\exp\left(\frac{\ln N_{\max}-\ln N_{\min}}{L-1}\right)$
$N_{max}$is chosen to match the finest detail in the training data. We use spatial hash function:
$h(\mathbf{x})=\left(\bigoplus_{i=1}^{d}x_{i}\pi_{i}\right)\quad\bmod T$ (4)
where $\bigoplus$ denotes the bit-wise XOR operation and $\pi_{3}$ are large prime numbers. Effectively, this formulation applies an XOR operation to the outputs of per dimension linear congruential pseudo random permutations, thereby decorrelating the influence of individual dimensions on the resulting hash value. We choose $\pi_{1}:=1$ for better cache coherence, $\pi_{2}=2654435761$, and $\pi_{3}=805459861$.
Then the background decoder maps a coordinate $x$, direction $d$ and scene feature to its density and color:
${[\sigma(\boldsymbol{x}),\boldsymbol{z}]}$ $=F_{\theta_{bkg,1}}\left(Hashenc(\boldsymbol{x})\right)$ (5)
@@
Background Model Architecture For the background model, our first stage uses 8 fully-connected layers and the second stage uses 4 fully-connected layers. All layers have a hidden size of 256. Further architecture details are shown in Table. 1. The background representation model is implemented as a two staged multilayer perceptron. The inputs to the background model are exactly the same as NeRF. The inputs consist of a 3D point given in Cartesian coordinates $x=x,y,z$ to the first stage and a 3D direction with $d=d_{x},d_{y},d_{z}$ as input to the second. We use a Fourier feature mapping, which increases training speed for low dimensional features. The 3D location x is mapped with a Fourier encoding at k = 10 frequencies for
$\gamma(p)=[...,\sin(2^{k}\pi p),\cos(2^{k}\pi p),...]\text{ for }k=0,...,K-1$
resulting in a length 63 input vector to the first stage. 8 fully-connected layers form the first stage, that outputs a density value and an intermediate feature vector of size 256. The Fourier encoding for the direction uses 4 frequencies and the encoded vector is concatenated with the output feature vector y(x) of size 256, resulting in a input vector of length 283 to the second stage. The 4 layer deep second stage outputs a color value.
3.3 Frequency Modulated Foundation Model
All existing NeRF methods have difficulties in multi-dynamic objects in urban scenes. The original NeRF method assumes the density $\sigma$ is static, but both the density $\sigma(\boldsymbol{x})$ and color $\boldsymbol{c}(\boldsymbol{x})$ depend on time (and video). SUDS [suds] and Emernerf [emernerf] use an optical flow method to differentiate dynamic objects from the background. However, their method cannot model dynamic objects effectively, resulting in blurred object portions and numerous artifacts in the rendered images. Furthermore, most of the dynamic objects only appear in a few observations. This setting starkly contrasts the majority of existing works that observe multiple views of the same object. To address this challenging problem, we propose a frequency auto-encoder in dynamic objects node to learn a normalized, object-centric high-level feature that describes and disentangles shape, appearance, and pose.
Auto-encoder is a type of neural network architecture used for unsupervised learning of efficient encodings of input data. They consist of an encoder network that maps the input to a latent code, and a decoder network that reconstructs the input from the code. We employ a vision foundation model as the encoder to extract accurate shape and appearance representations, benefiting from the strong priors learned through large scale pretraining. This design enables the proposed method to more effectively address the sparse observation challenges of dynamic objects in autonomous driving scenarios. The code is a compressed representation that encodes the most salient features of the data.
@@ 3.3 Frequency Modulated Foundation Model
where $S_{o}$ is a matrix with the inverse length values in the middle. $s_{o}=[L_{o},H_{o},W_{o}]$ with $S_{o}=diag(1/s_{o})$. The scaling factor allows the network to learn size-independent representation. We share the weights of the network for the objects of the same class (such as cars, vans and trucks that were categorized during preprocessing.)
@@
Shape and Appearance Encoder. For a input image $I$, we use the segment anything model and 3D object detector to get the 2D mask $M$ and 3D bounding box $\beta$. We encode the input images $I$ depicting a given object of interest into a shape code $\mathcal{S}_{\theta}$ and an appearance code $\mathcal{A}_{\theta}$ via a neural network. Inspired by [autorf], we design a two-head output architecture with the vision foundation extractor for generating the shape code $l_{s}$ and appearance code $l_{a}$, respectively.
We use the DINOv2 backbone to extract image features. The output tokens of the DINO model consist of both a class token and patch tokens. The class token primarily encodes high-level semantic information, which we map through a lightweight MLP to generate the shape code of the dynamic object. The patch tokens mainly capture local appearance cues such as color and texture. We apply a mean pooling operation over the patch tokens, followed by another lightweight MLP, to obtain the appearance code. The final latent representation comprises both the shape code and appearance code, each with 128 dimensions. Image inputs are scaled to a maximum of 320 pixels in each dimension, maintaining the original aspect ratio.
Table 2: Object Decoder Architecture.
Object Decoder Model
Layer Name Input Size
density layer0 Agg.Block+ResnetBlockFC 128128
density layer1 Agg.Block+ResnetBlockFC 128128
density layer2 Agg.Block+ResnetBlockFC 128128
density layer3 Agg.Block+ResnetBlockFC 128128
density layer4 Agg.Block+ResnetBlockFC 128128
density lin_out Linear 11
color layer0 Agg.Block+ResnetBlockFC 128128
color layer1 Agg.Block+ResnetBlockFC 128128
color layer2 Agg.(+density_feats)+ResnetBlockFC 128128
color layer3 Agg.(+query_cams)+ResnetBlockFC 128128
color layer4 Agg.(+query_cams)+ResnetBlockFC 128128
color lin_out Linear 33
Shape and Appearance Decoder $\phi$. The shape code $l_{s}$ is fed to the shape decoder network $\phi_{s}$ which implicitly outputs an density value $\sigma$ of the given coordinatexx. Unlike the shape code decoder, the appearance code decoder $\phi_{a}$ uses both shape and appearance codes as the input with the positional and directional information. The appearance decoder also takes the dynamic object node information to perceive the object’s position better and implicitly outputs the RGB color $c$. In Tab. 2, we present architecture of object decoder.
Shape Decoder. The shape decoder is a MLP that is made of 5 ResNet blocks with hidden dimension 128. At each layer, we feed the previous feature map and the positional encoding of the query points. In order to match the dimensionality of the positional encoding with the hidden dimensions of the MLP, we apply a single linear layer and aggregate the output with the intermediate feature maps by a simple per-channel mean pooling.
Color Decoder. For decoding the color, we use a similar architecture as for the shape decoder: A MLP of 5 ResNet blocks with hidden dimensionality of 128 and additional linear aggregations for additional input: As for the shape decoder, we aggregate intermediate features with positional encodings by mean pooling. Furthermore, on the third layer we pass in the same way the output of the corresponding layer of the shape encoder. This enables the color decoder to incorporate estimated shape information. On the final two layers, we pass the view direction (encoded as 3-dimensional vector) to account for view-dependent effects.
@@
Frequency Modulation. In real-world urban scenes, dynamic objects only appear in a few observations, which significantly differs from the settings of the original NeRF method. In sparse view settings, the Neural radiance field is prone to overfitting to these 2D images without explaining 3D geometry in a multi-view consistent way, leading to reconstruction failure. Prior work [freenerf] has already demonstrated that this issue is presumably caused by the high-frequency feature of the input.
@@
To this end, we further improve the feature-extracting foundation model with a frequency regularization method. The original method addressed the sparsity issue by utilizing a CNN network to encode the high-frequency features, which leads to the overfitting of high-frequency features. In this work, we use a linearly increasing frequency mask $\boldsymbol{\alpha}$ to regulate the visible frequency spectrum of the position embedding function:
@@
Frequency regularization restricts the high-frequency components of the neural network weights or activations during training. This prevents the model from overfitting to noisy high-frequency patterns in the data. It acts as a low-pass filter, smoothing the weight updates and forcing the model to learn more general, low-frequency representations, ultimately improving generalization. All the additional information related to scene graph dynamic objects nodes: the coordinate $\boldsymbol{x}$, direction $\theta\phi$, and object pose is embedded by our frequency encoding before they are put into the auto-encoder.
Our method addresses the reconstruction of dynamic objects from sparse observations, distinct from prior approaches such as FreeNeRF [freenerf] that exclusively focus on static objects. Unlike FreeNeRF [freenerf], which encodes only input points, our frequency auto-encoder architecture incorporates dynamic object poses, input points, and appearance and shape features from scene graph nodes. This adaptation significantly enhances the model’s capacity to handle complex dynamic scenes, surpassing the capabilities of the previous method. Our frequency auto-encoder significantly improves the scene representation of dynamic objects and achieves great performance in image reconstruction, novel view synthesis, and object editing.
Figure 5: Qualitative results on reconstruction and novel scene arrangements of a scene from the KITTI dataset [kitti] for NeRF [NeRF], NSG [nsg], SUDS [suds], MARS [mars], PVG [pvg] and our method. From left to right, these images correspond to different timesteps captured in the dynamic scene. We place the PSNR values of each scene in the bottom right corner. NeRF and SUDS are limited in their ability to properly represent the dynamic parts of the scene. In contrast, our neural scene graph method achieves high-quality view synthesis results for both reconstruction and novel scene synthesis, regardless of scene dynamics.
Table 3: Image reconstruction and novel view synthesis results on the complete KITTI [kitti] dataset. We compare our method with others on PSNR, SSIM, and LPIPS metrics. For PSNR and SSIM, higher is better; for LPIPS, lower is better. Our method generates higher-quality rendering images both in image reconstruction and novel view synthesis.
@@
KITTI [kitti] KITTI - 75% KITTI - 50% KITTI - 25%
@@
NeRF [NeRF] 16.98 0.509 0.524 16.81 0.509 0.524 16.64 0.504 0.535 16.55 0.498 0.544
NeRF++Time 24.18 0.677 0.237 22.05 0.612 0.262 21.90 0.602 0.239 20.17 0.591 0.253
NSG [nsg] 26.66 0.857 0.071 25.25 0.801 0.089 24.81 0.786 0.105 22.31 0.752 0.134
SUDS [suds] 28.31 0.857 0.065 25.21 0.834 0.144 24.29 0.821 0.156 19.73 0.758 0.217
S-NeRF [snerf] 19.25 0.664 0.193 18.71 0.606 0.352 18.31 0.598 0.361 17.93 0.586 0.365
3DGS [3dgs] 21.02 0.811 0.202 19.54 0.776 0.224 18.95 0.763 0.232 18.27 0.758 0.236
Mars [mars] 27.96 0.900 0.105 24.31 0.845 0.102 24.19 0.601 0.118 22.85 0.732 0.125
EmerNeRF [emernerf] 29.40 0.921 0.061 26.32 0.862 0.089 25.02 0.851 0.092 22.98 0.819 0.122
StreetGaussian [streetgaussian] 29.45 0.926 0.059 26.79 0.844 0.083 25.31 0.841 0.089 24.54 0.828 0.098
PVG [pvg] 29.37 0.922 0.061 26.43 0.839 0.086 25.12 0.837 0.091 24.23 0.823 0.101
Ours 30.31 0.931 0.057 27.75 0.887 0.080 26.55 0.878 0.086 25.69 0.856 0.091
Table 4: We report the PSNR, SSIM, and LPIPS metrics results on VKITTI2 [vkitti] dataset. Our approach outperforms prior methods of reconstruction and novel view synthesis across all evaluation metrics.
Methods Year Image Reconstruction Novel View Synthesis
VKITTI2 [kitti] VKITTI2 - 75% VKITTI2 - 50% VKITTI2 - 25%
[…]
NeRF [NeRF] ECCV’20 17.63 0.525 0.511 17.11 0.513 0.514 16.94 0.517 0.521 16.65 0.504 0.537
NeRF + Time 24.54 0.683 0.235 22.59 0.619 0.252 22.15 0.612 0.261 21.82 0.615 0.233
NSG [nsg] CVPR’21 27.18 0.894 0.092 25.45 0.886 0.094 24.95 0.794 0.101 23.49 0.758 0.105
SUDS [suds] CVPR’23 28.34 0.887 0.049 27.52 0.864 0.064 26.25 0.845 0.074 21.47 0.788 0.088
S-NeRF [snerf] ICLR’23 20.53 0.874 0.181 19.46 0.646 0.342 19.11 0.604 0.354 18.43 0.591 0.359
StreetSurf [streetsurf] Arxiv’23 24.89 0.839 0.242 23.88 0.785 0.273 22.79 0.776 0.289 21.59 0.746 0.303
3DGS [3dgs] TOG’24 22.34 0.868 0.179 21.42 0.657 0.339 21.11 0.624 0.347 20.84 0.863 0.329
Mars [mars] CICAI’23 30.26 0.925 0.083 29.79 0.917 0.088 29.63 0.916 0.087 27.01 0.877 0.104
Streetgaussian [streetgaussian] ECCV’24 31.37 0.928 0.081 30.08 0.922 0.084 30.03 0.910 0.089 28.13 0.905 0.099
PVG [pvg] IJCV’26 31.69 0.931 0.077 30.45 0.925 0.080 30.15 0.914 0.086 28.45 0.910 0.095
Ours 32.42 0.938 0.075 31.58 0.935 0.079 30.95 0.931 0.084 29.69 0.927 0.088
@@ 3.4 Scene Graph Rendering
Ray-Plane Intersection  Ray-Plane Intersection focuses on efficiently modeling static components using sparse planes. These planes are uniformly distributed along the scene’s depth, intersecting rays to generate sampling points for the static background. In our experiment we use 10 to 32 planes to capture different movement scenarios and camera perspectives.
Ray-Box Sampling. For each ray $r$ in the dynamic object node, we use the ray-box sampling method. We check the rays that intersect with the dynamic object nodes and translate them to the object coordinate space. Then, we apply the efficient AABB intersection method [aabb] to compute all ray-box entrance and exit points $(t_{o1},t_{oN})$. We stratified sample points $\boldsymbol{x}_{o}$ between the entrance and exit points:
@@
Ray-Box Intersection involves transforming rays into local frames of dynamic objects, enabling computation of intersections within scaled axis-aligned bounding boxes (AABBs), as shown in Fig. 4. Equidistant quadrature points along the rays within the bounding box are sampled to represent object surfaces accurately. Experimentally, using 7 equidistant points per ray and object node was found to balance accuracy and rendering speed.
[…]
@@
Figure 7: We also present our method for novel view synthesis performance on the KITTI 0000 sequence, with high-dynamic human. The results clearly show that our approach achieve high accuracy reconstruction perfromance.
Figure 9: We also present our method for scene reconstruction performance. The results clearly show that our approach achieve high accuracy reconstruction perfromance.
Figure 10: We also present our method for scene reconstruction performance on the Waymo dataset [waymo], compared with PVG and Desire-gs. The results clearly show that our approach achieve high accuracy reconstruction perfromance.
Table 5: Reconstruction performance on Waymo [waymo] dataset. We outperform other works on image reconstruction and novel view synthesis.
Methods Waymo [waymo]
Image Reconstruction Novel View Synthesis
[…]
NeRF [NeRF] 18.89 0.515 0.398 18.45 0.508 0.405
NSG [nsg] 24.08 0.656 0.441 21.05 0.575 0.487
SUDS [suds] 28.83 0.805 0.317 25.36 0.783 0.384
S-NeRF [snerf] 19.67 0.528 0.387 19.25 0.517 0.402
StreetSurf [streetsurf] 26.73 0.849 0.372 23.79 0.825 0.401
3DGS [3dgs] 27.99 0.866 0.293 25.08 0.822 0.319
MARS [mars] 21.84 0.682 0.431 20.69 0.637 0.453
EmerNeRF [emernerf] 28.11 0.784 0.375 25.94 0.765 0.383
PVG [pvg] 31.45 0.897 0.243 27.49 0.838 0.303
DeSiRe-GS [desiregs] 31.89 0.904 0.238 28.84 0.842 0.300
Ours 32.55 0.912 0.235 29.89 0.861 0.275
Table 6: Quantitative comparison with existing methods on the nuScenes dataset [nuScenes]. We report PSNR, SSIM, and LPIPS. Higher PSNR and SSIM indicate better performance, while lower LPIPS is better.
[…]
Instant-NGP [instantngp] 16.78 0.519 0.570
NeRF+Time 17.54 0.565 0.532
NSG [nsg] 21.67 0.671 0.424
Mip-NeRF [mipnerf] 18.08 0.572 0.551
Mip-NeRF360 [mipnerf360] 22.61 0.688 0.395
Urban-NeRF [urbannerf] 20.75 0.627 0.480
S-NeRF [snerf] 25.43 0.730 0.302
SUDS [suds] 21.26 0.603 0.466
EmerNeRF [emernerf] 26.75 0.760 0.311
3DGS [3dgs] 26.08 0.717 0.298
4DGS [4dgs] 19.79 0.622 0.473
DrivingGaussian [drivinggaussian] 28.74 0.865 0.237
Ours 29.24 0.878 0.223
@@ 3.5 Progressive Scene Graph Joint Learning
For the large-scale dynamic scene, we divide the entire scene into multi-local scene graphs. Whenever the camera moves to the bound of the current scene graph, we dynamically initialize a new neural scene graph and stop optimizing the previous ones.
When encountering the same objects observed in two local scenes, we update only the dynamic object network in the current scene graph because the optimization of the previous scene graphs is frozen. To mitigate potential quality reduction in the former scene graph caused by this strategy, each subset contains some overlapping images, essential for achieving consistent reconstructions in the global scene. We also dynamically adjust local scene bounds during optimization. This helps prevent the undesired splitting of dynamic elements and optimizes the overall reconstruction process.
In each local scene graph, we use color loss to optimize the representation model for the background and dynamic object node:
@@
where $\Pi()$ is the projection from 3D point $P_{i}$ to the depth of 2D pixel. $R_{i}$ is the set of rays with a valid depth projection value. Inspired by [dnerf], we also use ray distribution loss to minimize the KL divergence between the rendered ray distribution $h(t)$ and the noisy depth distribution:
$\mathcal{L}_{\sigma}=\sum_{i}\log{h_{i}}\mathrm{exp}(-\frac{(t_{i}-\hat{\boldsymbol{D}_{i}})^{2}}{2\hat{\sigma}_{i}^{2}})\delta_{i}$ (14)
@@
$\mathcal{L}_{\mathrm{seg}}=\mathbb{E}_{\mathbf{r}\sim I_{i}}\left[\mathcal{S}_{i}(\mathbf{r})\int_{t_{n}}^{t_{f}}(T\cdot\alpha)^{2}dt\right]$ (15)
@@
4 Implementation
4.1 Dataset and Metrics
We evaluate ProSGNeRF on KITTI tracking datasets [kitti] for the real-world urban scene. The KITTI dataset [kitti] is captured using a stereo camera. Each sequence consists of more than 100 time steps and images of size $1242\times 375$, each from two camera perspectives and more than 12 dynamic objects from different classes. We strictly follow the experimental setting of NSG and SUDS to ensure fairness. In addition to the experiments on the KITTI dataset, we further conducted evaluations on the synthetic vKITTI2 dataset [vkitti] and the larger real-world Waymo dataset [waymo] and nuScenes [nuScenes] datasets, thereby demonstrating the outstanding performance of our model across diverse scenarios and ensuring a comprehensive evaluation. We report the standard view synthesis metrics PSNR, SSIM, and LPIPS scores for all evaluations.
Dataset Generation We construct a new dataset using the pre-trained SAM2 model and a 3D object detection model, which contains camera intrinsics and extrinsics, RGB images, as well as 3D bounding boxes and 2D masks of objects. We use SAM2 to acquire precise 2D semantic masks without the requirement for human annotation. Leveraging the promptable segmentation method, SAM predicts masks for both dynamic objects and scenarios in far-field situations. This approach enables the construction of a progressive neural scene graph, integrating instance masks from SAM and 3D bounding boxes derived from machine predictions.
4.2 Parameters
ProSGNeRF is evaluated on a desktop PC with NVIDIA A100 GPU. For experimental settings. Each decoder uses the five fully-connected residual blocks. The bounding box dimensions are scaled to include the shadows for better reconstruction.
Loss weight  In refining local scene graphs, we utilize a color loss mechanism to enhance the representation of background and dynamic object nodes. Introducing a novel lidar points projection loss bolsters geometric consistency by supervising rendering depth. Drawing inspiration from [dnerf], our approach incorporates a ray distribution loss to minimize divergence between the rendered ray distribution $h(t)$ and observed noisy depth distribution.Moreover, we introduce a far-field loss function to ensure zero density for point samples along rays passing through sky pixels. The depth loss and KL divergence loss are pivotal in maintaining geometric consistency, ultimately enhancing scene accuracy and geometry portrayal. However, excessive emphasis on these losses can negatively impact image metrics. Hence, controlling their weights is crucial for optimal performance.
In our experiments, We set $\lambda_{c}=1$ , $\lambda_{d}=0.005$ $\lambda_{\sigma}=0.005$, $\lambda_{seg}=0.001$. These adjustments strike a balance between geometric accuracy and image fidelity, refining scene representation while minimizing adverse effects on image-related metrics.
Baseline We compare our method against state-of-art implicit neural rendering methods: the original NeRF implementation [NeRF], NeRF+Time, NSG [nsg], SUDS [suds], S-NeRF [snerf], StreetSurf [streetsurf], MARS [mars], EmerNeRF [emernerf], and 3D Gaussian Splatting [3dgs], StreetGaussian [streetgaussian], Driving Gaussian [drivinggaussian], PVG [pvg], DeSiRe-GS [desiregs].